# Dual Collaborative LLMs via Continual Fine-Tuning for Serendipitous Recommendation

### v2

@@ 1. Introduction
Large language models (LLMs) (Achiam et al., 2023; Guo et al., 2025; Yang et al., 2025; Chen et al., 2025), with their powerful reasoning capabilities and massive knowledge, can generate diverse recommended content. Existing methods (Chen, 2021; Lin et al., 2024; Wang et al., 2024) reduce reliance on historical behaviors by introducing novel content to enhance long-term engagement. As shown in Figure 1, current methods (Wang et al., 2025a, b; Li et al., 2023) adopt a hierarchical framework: using the Novelty LLM to explore diversity and the Relevance LLM to handle relevance, respectively. However, two key challenges remain:
@@ 2.1. LLMs for Recommendation
Current scholarly efforts can be broadly classified into two principal directions. The first strategy capitalizes on LLMs’ intrinsic aptitude for text processing to fulfill recommendation tasks (Harte et al., 2023; He et al., 2023; Hou et al., 2024; Gao et al., 2023; Dai et al., 2023). It has been observed that LLMs demonstrate a noteworthy ability to make recommendations through straightforward in-context learning. This finding indicates that the models, even during pretraining, can discern item relationships and latch onto fundamental collaborative signals, thus inspiring further research into the integration of LLMs within recommendation systems. The second avenue leverages LLMs for data augmentation (Wei et al., 2024; Liu et al., 2024) and representation learning (Ren et al., 2024), aiming to bolster traditional recommendation systems. In addition to optimizing basic recommendation capabilities, LLMs also provide a new perspective for solving the Exploration-and-Exploitation (EE) balance problem (Auer et al., 2002; Coppolillo et al., 2024; Li et al., 2023; Vargas and Castells, 2011; Zhao et al., 2024; Bianchi, 2025) in recommendation systems. EXPLORE (Coppolillo et al., 2024) constructs a probabilistic user behavior model and integrates relevance with diversity via a copula function to develop a novel recommendation strategy, aiming to maximize the total amount of knowledge users are exposed to. Google-v1 (Wang et al., 2025a) combines hierarchical planning with LLM inference scaling, and separates the goals of novelty and user alignment. Google-v2 (Wang et al., 2025b) further introduces multimodal large models on this basis, discovers users’ new interests through the chain-of-thought strategy. However, these methods still have limitations: they fail to fully utilize the user group preference information contained in users’ long-term interests.
@@ 4.2. Dual-Stable Interest Exploration
Existing serendipitous recommendation method (Wang et al., 2025a) mines novel items that users may be interested in by modeling their short-term behavior sequences. However, such methods fail to fully capture the long-term stable preferences implied by users’ group identities from short-term sequences. To address this issue, we propose the Dual-Stable Interest Exploration (DSIE) module
@@ 4.3. Periodic Collaborative Optimization
Existing methods (Wang et al., 2025a), after obtaining feedback from the Relevance LLM and completing the alignment of the Novelty LLM, only generate a static offline database. This makes them unable to continuously integrate dynamically updated user behavior information and also causes the system to fail to achieve closed-loop optimization based on incremental data. To address this issue, we propose a Periodic Collaborative Optimization (PCO) mechanism: by regularly feeding back user preferences learned by the Relevance LLM to the Novelty LLM for incremental alignment, and providing the output of the incrementally fine-tuned Novelty LLM to the Relevance LLM for re-scoring, thereby constructing a sustainably evolving dual-model iterative loop.
@@ 5.2. Overall Performance Comparison (RQ1)
We compare our CoEA with 4 exploitation-based baseline models (i.e., Two-tower (Yang et al., 2020), GRU4Rec (Hidasi et al., 2015), SASRec (Kang and McAuley, 2018), and Bert4Rec (Sun et al., 2019)) and 7 exploration-based baseline models (i.e., HUCB/pHUCB (Song et al., 2022), NLB (Su et al., 2024), LLM-KERec (Zhao et al., 2024), EXPLORE (Coppolillo et al., 2024), Google-v1 (Wang et al., 2025a), and Google-v2 (Wang et al., 2025b)). Since the two datasets do not contain multimodal information, we use a text LLM instead of a multimodal LLM when conducting experiments on the Google-v2 model. Details of the baseline models are available in Appendix A.5, and implementation details of CoEA are in Appendix A.6. The experimental results are the average value obtained after running five times. Table 2 presents the overall performance on the two datasets, with the result analysis as follows:
@@ A.5. Baselines
To verify the effectiveness of our method, we compare it with representative baseline models, including 4 exploitation-based baselines: Two-tower (Yang et al., 2020), GRU4Rec (Hidasi et al., 2015), SASRec (Kang and McAuley, 2018), and Bert4Rec (Sun et al., 2019), as well as 7 exploration-based baselines, namely HUCB (Song et al., 2022), pHUCB (Song et al., 2022), NLB (Su et al., 2024), LLM-KERec (Zhao et al., 2024), EXPLORE (Coppolillo et al., 2024), Google-v1 (Wang et al., 2025a), and Google-v2 (Wang et al., 2025b).
@@
  • Google-v1 (Wang et al., 2025a): An approach that combines hierarchical planning with LLM inference scaling and distinguishes between the goals of novelty and user alignment.
  • Google-v2 (Wang et al., 2025b): A framework that introduces multimodal large models based on Google-v1 and discovers users’ new interests through the chain-of-thought strategy.

### v3

# When Relevance Meets Novelty: Dual-Stable Periodic Optimization for Serendipitous Recommendation

@@ 1. Introduction
Large language models (LLMs) (Achiam et al., 2023; Guo et al., 2025; Yang et al., 2025; Chen et al., 2025), with their powerful reasoning capabilities and massive knowledge, can generate diverse recommended content. Existing methods (Chen, 2021; Lin et al., 2024; Wang et al., 2024) reduce reliance on historical behaviors by introducing novel content to enhance long-term engagement. As shown in Figure 1, current methods (Wang et al., 2025b, a; Li et al., 2023) adopt a hierarchical framework: using the Novelty LLM to explore diversity and the Relevance LLM to handle relevance, respectively. However, two key challenges remain:
@@ 2.1. LLMs for Recommendation
Current scholarly efforts can be broadly classified into two principal directions. The first strategy capitalizes on LLMs’ intrinsic aptitude for text processing to fulfill recommendation tasks (Harte et al., 2023; He et al., 2023; Hou et al., 2024; Gao et al., 2023; Dai et al., 2023). It has been observed that LLMs demonstrate a noteworthy ability to make recommendations through straightforward in-context learning. This finding indicates that the models, even during pretraining, can discern item relationships and latch onto fundamental collaborative signals, thus inspiring further research into the integration of LLMs within recommendation systems. The second avenue leverages LLMs for data augmentation (Wei et al., 2024; Liu et al., 2024) and representation learning (Ren et al., 2024), aiming to bolster traditional recommendation systems. In addition to optimizing basic recommendation capabilities, LLMs also provide a new perspective for solving the Exploration-and-Exploitation (EE) balance problem (Auer et al., 2002; Coppolillo et al., 2024; Li et al., 2023; Vargas and Castells, 2011; Zhao et al., 2024; Bianchi, 2025) in recommendation systems. EXPLORE (Coppolillo et al., 2024) constructs a probabilistic user behavior model and integrates relevance with diversity via a copula function to develop a novel recommendation strategy, aiming to maximize the total amount of knowledge users are exposed to. Google-v1 (Wang et al., 2025b) combines hierarchical planning with LLM inference scaling, and separates the goals of novelty and user alignment. Google-v2 (Wang et al., 2025a) further introduces multimodal large models on this basis, discovers users’ new interests through the chain-of-thought strategy. However, these methods still have limitations: they fail to fully utilize the user group preference information contained in users’ long-term interests.
@@ 4.2. Dual-Stable Interest Exploration
Existing serendipitous recommendation method (Wang et al., 2025b) mines novel items that users may be interested in by modeling their short-term behavior sequences. However, such methods fail to fully capture the long-term stable preferences implied by users’ group identities from short-term sequences. To address this issue, we propose the Dual-Stable Interest Exploration (DSIE) module
@@ 4.3. Periodic Collaborative Optimization
Existing methods (Wang et al., 2025b), after obtaining feedback from the Relevance LLM and completing the alignment of the Novelty LLM, only generate a static offline database. This makes them unable to continuously integrate dynamically updated user behavior information and also causes the system to fail to achieve closed-loop optimization based on incremental data. To address this issue, we propose a Periodic Collaborative Optimization (PCO) mechanism: by regularly feeding back user preferences learned by the Relevance LLM to the Novelty LLM for incremental alignment, and providing the output of the incrementally fine-tuned Novelty LLM to the Relevance LLM for re-scoring, thereby constructing a sustainably evolving dual-model iterative loop.
@@ 5.2. Overall Performance Comparison (RQ1)
We compare our CoEA with 4 exploitation-based baseline models (i.e., Two-tower (Yang et al., 2020), GRU4Rec (Hidasi et al., 2015), SASRec (Kang and McAuley, 2018), and Bert4Rec (Sun et al., 2019)) and 7 exploration-based baseline models (i.e., HUCB/pHUCB (Song et al., 2022), NLB (Su et al., 2024), LLM-KERec (Zhao et al., 2024), EXPLORE (Coppolillo et al., 2024), Google-v1 (Wang et al., 2025b), and Google-v2 (Wang et al., 2025a)). Since the two datasets do not contain multimodal information, we use a text LLM instead of a multimodal LLM when conducting experiments on the Google-v2 model. Details of the baseline models are available in Appendix A.5, and implementation details of CoEA are in Appendix A.6. The experimental results are the average value obtained after running five times. Table 2 presents the overall performance on the two datasets, with the result analysis as follows:
@@ A.5. Baselines
To verify the effectiveness of our method, we compare it with representative baseline models, including 4 exploitation-based baselines: Two-tower (Yang et al., 2020), GRU4Rec (Hidasi et al., 2015), SASRec (Kang and McAuley, 2018), and Bert4Rec (Sun et al., 2019), as well as 7 exploration-based baselines, namely HUCB (Song et al., 2022), pHUCB (Song et al., 2022), NLB (Su et al., 2024), LLM-KERec (Zhao et al., 2024), EXPLORE (Coppolillo et al., 2024), Google-v1 (Wang et al., 2025b), and Google-v2 (Wang et al., 2025a).
@@
  • Google-v1 (Wang et al., 2025b): An approach that combines hierarchical planning with LLM inference scaling and distinguishes between the goals of novelty and user alignment.
  • Google-v2 (Wang et al., 2025a): A framework that introduces multimodal large models based on Google-v1 and discovers users’ new interests through the chain-of-thought strategy.